OpenStudy (anonymous):

Using L'Hospital's rule, solve lim x->0 of e^(x)-1-x/x^(2)

OpenStudy (anonymous):

Is all of the text before the "/" the numerator?

OpenStudy (anonymous):

Assuming it is, e^(0)-1 = 1 - 1 =0, it is lim x->0 of -x/x^2, this ends up being 0/0, so we use L'Hospital's again (take derivative of top and bottom individually and take limit again)--> lim x->0 of -1/(2*x) = -1/0 = -infinity

OpenStudy (anonymous):

if this I see it right(use L'H rule twice): \[\lim (e ^{x}-1-x)/x ^{2}=\lim (e ^{x}-1)/2x=\lim e ^{x}/2=1/2\]

OpenStudy (anonymous):

Oops, inik is correct

OpenStudy (anonymous):

sorry...i typed it in wrong myininaya. the stuff before the "/" is the numerator, and the x^2 is the denominator.

myininaya (myininaya):

ok

myininaya (myininaya):

what inik has is corrrect then